OpenStudy (anonymous):

Solve 5x2 = -30x - 65 (2 points)

OpenStudy (campbell_st):

divide every term by 5 and rewrite the equation \[x^2 +6x + 13 = 0\] now use the general quadratic formula or completing the square

OpenStudy (campbell_st):

your problem has complex roots...

OpenStudy (uri):

5x2 = -30x - 65 5x^2+30x-65=0 divide it by 5 So, x^2+6x-13=0 Use Quadratic formula: